OpenStudy (anonymous):

Write as a single trig function or a power of a single trig function. Cot(theta)/Tan(theta)

terenzreignz (terenzreignz):

It must have somethng to do with the fact that \[\huge \frac1{\tan (\theta)}=\cot(\theta)\]

OpenStudy (ajprincess):

\(\tan\theta=1/\cot\theta\) Does that help?

OpenStudy (dan815):

oh cot

terenzreignz (terenzreignz):

Yeah, that's what I thought first, too, @dan815 :)

OpenStudy (dan815):